OpenStudy (anonymous):

PLZ HELP THIS GUY MY GOD PLZ y = csc(1/2x) ... i, trying to figure out the period change BEEN HERE FOR A WHILE NOW

OpenStudy (sweetburger):

its not csc(theta/2)?

OpenStudy (anonymous):

nope

OpenStudy (sweetburger):

so i believe the period would be 1pi

OpenStudy (anonymous):

csc=1/sin csc has the same period as sin

OpenStudy (anonymous):

the period of sin is 2pi. the period of sin(x/2) is 4pi.

OpenStudy (anonymous):

since sin(x+2pi)=sin(x) and sin((x+4pi)/2)=sin(x/2)

OpenStudy (anonymous):

so like pi/1/2 = 4pi | 4pi/2 = 2/pi | pi/2/2 = pi ???